a grain retailer puts a 100kg lead block under the platform of his scale. by how much does this increase the reading of the scale when 50 kg sack of rice is on the top of its platform, if the distance between their centers are .03m apart


1
Expert's answer
2022-01-13T13:25:00-0500

"F=\frac{Gm_1m_2}{r^2}=0.37~mN,"

"\Delta m=\frac Fg=0.038~g."